Using VSEPR Theory, name and sketch the molecular shape of each of the molecules listed:
a) C2Br4(s)
b) HgCl2(s)
c) SnCl4(aq)
1
Expert's answer
2019-08-08T04:43:07-0400

1) C2Br4

Valence electrons = 4*2+7*4 =36

Lewis dot structure



For each carbon atom:

bonding pairs =3

nonbonding pairs =0

general formula AB3

molecular geometry - trigonal planar

bond angle =120


2) HgCl2

Valence electrons 2+7*2 = 16

Lewis dot structure



bonding pairs = 2

nonbonding pairs = 0

general formula AB2

molecular geometry = linear

bond angle =180


3) SnCl4

Valence electrons = 4+7*4= 32

Lewis dot structure



bonding pairs = 4

nonbonding pairs =0

general formula = AB4

molecular geometry - tetrahedral

bond angle = 109.5